# Multi-color characterization of optically invisible FU Orionis-type outbursts: Demonstration and prospects for the WINTER survey

### v1

@@ Abstract
1 Introduction
Young stellar objects (YSOs) grow through episodic accretion, where material from a surrounding disk falls onto the star in bursts rather than at a steady rate (Fischer et al., 2023, and references therein). These accretion outbursts significantly increase the system’s luminosity and play a major role in stellar mass assembly (Kenyon et al., 1990; Hartmann & Kenyon, 1996; Vorobyov & Basu, 2005). Episodic accretion has been proposed as a solution to the “luminosity problem,” in which protostars appear dimmer than expected if they were gaining mass at a constant rate throughout their evolution (Kenyon et al., 1990; Dunham & Vorobyov, 2012). In addition to influencing how stars gain mass, these bursts can alter the thermal and chemical structure of the disk, affecting processes such as snowline migration and early planet formation (Nayakshin et al., 2023; Lee et al., 2019; Jørgensen et al., 2013; Kóspál et al., 2023). Discovering and characterizing these outbursts remains essential for understanding their physical origins and role in star formation.
The most extreme form of episodic accretion in YSOs is FU Orionis-type (FUor) outbursts, characterized by large luminosity increases (mag 4–6) and lasting decades or more (Hartmann & Kenyon, 1996; Audard et al., 2014; Fischer et al., 2023). FUor outbursts sustain accretion rates of up to , contributing significantly to stellar mass growth in the earliest stages of evolution (Hartmann & Kenyon, 1996; Vorobyov & Basu, 2015). In contrast, another common category of bursts, EX Lupi-type outbursts (EXors), are less luminous (mag 2–3), shorter-lived (weeks to a few years), and contribute less significantly to the stellar mass budget (Ashraf et al., 2024; Cruz-Sáenz de Miera et al., 2023). Despite the proposed significance of FUors in mass accumulation, they are observationally rare. The classic compilation of Connelley & Reipurth (2018) reported fewer than 35 FUor-type events had been confirmed using spectroscopic techniques, albeit with most objects having sparse or no photometric coverage of the outburst.
@@ 1 Introduction
Recent wide-field IR surveys have proven essential for uncovering outbursts in heavily embedded YSOs, including classes of FUor outbursts previously missed in optical monitoring. The VISTA Variables in the Via Lactea (VVV; Minniti et al. 2010) and its extension VVVX enabled the discovery of dozens of high-amplitude eruptive YSOs, including several FUor candidates in the Galactic plane (Contreras Peña et al., 2017a, b, 2024; Guo et al., 2024a). Similarly, mid-infrared (MIR) monitoring with WISE (Wright et al., 2010) and NEOWISE (Mainzer et al., 2014) uncovered a growing population of deeply embedded eruptive variables (Hillenbrand et al., 2018; Ashraf et al., 2024; Tran et al., 2024; Contreras Peña et al., 2025). While the VVV and NEOWISE searches were largely implemented via archival photometric searches together with spectroscopic follow-up many years after the outburst onset, the advent of real-time transient discovery techniques is enabling prompt identification and early spectroscopic follow-up of these outbursts. The ongoing Palomar Gattini-IR (PGIR) survey has uncovered several new NIR-bright FUor events in the northern hemisphere (Hillenbrand et al., 2021, 2025) and continues to provide a key source of detections, especially toward the Galactic plane (Earley et al., in prep.).
@@
In addition to discovery, multi-color infrared photometry and broad-band spectroscopy play a key role in confirming YSO outbursts and enabling detailed studies of accretion disk evolution. These observations also identify and exclude interlopers, including Mira variables, cataclysmic variables, and classical novae. Multi-wavelength monitoring probes accretion disk formation via changes in NIR and MIR colors—where delays between shorter and longer wavelength bands are critical to understanding the outburst triggering mechanism and instability propagation (outside-in or inside-out; Bell et al. 1995; Cleaver et al. 2023; Nayakshin & Elbakyan 2024). Simultaneously, such coverage traces the development of hot inner disk regions, the clearing or rebuilding of circumstellar material and the presence of infalling envelopes or outflows (Hillenbrand & Rodriguez, 2022; Caratti o Garatti et al., 2017). Connelley & Reipurth (2018) demonstrate that NIR spectroscopy provides a robust basis for classifying FUor-type objects, which exhibit a distinctive combination of strong molecular absorption, weak or absent emission lines, low-gravity spectral features, and a mixed-temperature spectrum that becomes progressively cooler at longer wavelengths.
@@
The Wide-Field Infrared Transient Explorer (WINTER) is a new NIR time-domain survey on a dedicated 1-meter robotic telescope at Palomar Observatory (Lourie et al., 2020; Frostig et al., 2024). Each pointing covers a 1.2 field of view with a 1.1” pixel scale and the option of observing in Y, J, or a shortened H (Hs) filters, centered on 1.02, 1.25, and 1.60 . WINTER covers the Galactic plane during the summer months, running multiple surveys at an approximately two-week cadence. Unlike VVV and PGIR which almost entirely operated their time-domain surveys in a single filter (in Ks and J band, respectively), WINTER carries out multi-color cadenced surveys of the Galactic plane. Like PGIR, WINTER also includes the infrastructure for rapid discovery of transients with a real-time image subtraction pipeline, operating with smaller field of view (1.2 vs. 25 ) but substantially higher depth (median vs. AB mag) and spatial resolution (1.1″vs 8.7″) than PGIR (De et al., 2020).
@@
We present the observations of WNTR24-cua and WNTR24-egv in WINTER, WISE, and other surveys, the source environment, and spectroscopic follow-up in Section 2. In Section 3, we analyze the observations, presenting a study of the host environments, the progenitors of the outbursts, and spectroscopically classify the FUors. We discuss the implications for future observations of YSO outbursts with WINTER and conclude in Section 4.
2 Observations
@@
2.1 Discovery, NEOWISE, and WINTER initial observations
@@ 2.2 Optical Observations
2.3 Light curve evolution
@@
WNTR24-cua was not detected in WISE until early 2016, when it began a sharp rise in the band, brightening by 1.7 magnitudes over 364 days (Figure 2). This was followed by a decline of 0.5 magnitudes in 2019, after which the source exhibited a more gradual increase of 0.62 magnitudes between 2020 and 2024. The initial rise was not seen in , likely due to the band’s 1 magnitude shallower depth; however, near peak brightness the source was detected at mag, with a color of mag. By 2024, the emission became bluer with mag. The brightening is also evident in VVV , which shows the onset of the outburst in 2016, with the progenitor detected prior to the rise in , , and . In contrast, Gaia -band observations do not detect the source up through May 2017, but then observe the outburst three months later, implying a rapid optical rise after the start of the NIR and MIR outburst.
@@
2.4 Archival imaging of the source environment
@@ 3.1 Environment
WNTR24-egv lies in close proximity to the H ii region [DLW84] G35.2S, with G35.2N also nearby.. Using Gaia EDR3 parallax measurements ( mas; Gaia Collaboration 2020), we estimate a distance of kpc to the region. The projected angular separation between WNTR24-egv and the center of G35.2S is , corresponding to a physical offset of only pc at this distance, indicating a likely association. A dense molecular core (W48 89) identified by Keown et al. (2019) is essentially identical in position with WNTR24-egv. Many sources in the immediate vicinity were identified by Kuhn et al. (2021) as YSOs, although not the progenitor of WNTR24-egv.
@@
Following the procedure outlined in Tran et al. 2024, we also estimate the distance to the outbursts by identifying nearby candidate YSOs with known distances. First, we take sources within 20’ of the FUors from the Spitzer (SEIP) source list (SSTSL2, Spitzer Science Center & Infrared Science Archive (IRSA)(2021), SSC) and the 2MASS All Sky Catalog of point sources (Cutri et al., 2003) and apply a color cut of and to identify sources with IR-color excess as likely YSOs. IR-excess contaminants, such as star-forming galaxies, are eliminated following the procedures outlined in Gutermuth et al. (2009).
@@
This results in 466 and 273 candidate YSOs for WNTR24-cua and WNTR24-egv respectively. For WNTR24-cua, 17 of the candidate YSOs have counterparts in the Gaia DR3 catalog (Gaia Collaboration, 2020) with measured parallaxes, yielding a median distance of kpc. Due to the discrepancies in the estimated distances presented in Guo et al. 2024a, we use the estimate of 2.4 kpc for the analyses in Section 3.3. Similarly, for WNTR24-egv, 10 of the candidate YSOs have Gaia distances ranging from 0.2 to 2.2 kpc. However, their relatively uniform spatial distribution offers no clear basis for a reliable distance estimate, and many of the stars in this region are likely too obscured for Gaia detections. We therefore adopt the distance of the nearby H ii region at 2.52 kpc.
3.2 Spectroscopic features and classification
@@
Both WNTR24-cua and WNTR24-egv are spectroscopically similar to other FUors at NIR wavelengths (as explored in detail in Connelley & Reipurth, 2018). The spectra are shown in Figure 3 with another FUor, Gaia17bpi (Hillenbrand et al., 2018), for comparison. All three spectra demonstrate a “mixed-temperature” spectrum characteristic of FUors. At the shortest infrared wavelengths (around 1 m), the spectra resemble a late-F to early-G type star at T6000 K. However, at longer wavelengths the spectra are consistent with an M-type star at T3000 K with deep molecular absorption features. For example, there is prominent absorption from the 12CO band-head (around ) and H2O band (m and m). Molecular absorption from VO and TiO is strongest in Gaia17bpi, present but weaker in WNTR24-cua, and not confidently detected in WNTR24-egv. Additionally, there are atomic lines from Al I, Ca I, Mg I, Na I, and Si I, along with H Paschen lines in both objects.
@@
Both of the outbursts show strong He I 10830 blue-shifted absorption, which is evidence for a strong wind (Connelley & Reipurth, 2018). We estimate a FWHM velocity of approximately 231 km/s and 305 km/s for WNTR24-cua and WNTR24-egv, respectively, and corrected for instrumental broadening. Measuring against the blue end of the absorption feature, we estimate a maximum velocity of 221 km/s and 321 km/s, for WNTR24-cua and WNTR24-egv.
@@
Based on these spectroscopic properties, we classify WNTR24-cua and WNTR24-egv as FUors following the criteria outlined in Connelley & Reipurth (2018). We measure the equivalent widths of CO, Na I, and Ca I, as outlined in Messineo et al. (2021), and compare to bona fide FUors from the Connelley & Reipurth (2018) sample (Figure 4). Both WNTR24-cua and WNTR24-egv fall within the bona fide sample, with notably large CO equivalent widths. Additionally, both sources exhibit the eight NIR spectral characteristics of FUors outlined in Connelley & Reipurth (2018). These characteristics include CO absorption and molecular band absorption, metallic, hydrogen, and He I lines, and changing of spectral type with wavelength.
3.3 Spectral energy distribution
@@
To place the SED-based extinction estimates in context, we also derive values from spectral dereddening and Galactic dust maps. First, we estimate the visual extinction () toward each source by dereddening their near-infrared spectra using the Cardelli et al. (1989) extinction law () until their continuum slopes match that of Gaia 17bpi, which has a known extinction of mag. This comparison yields relative extinctions of mag and mag, corresponding to values of mag for WNTR24-cua and mag for WNTR24-egv. We then compared these values to predictions from the Bayestar 3D Galactic dust map (Green et al., 2019). For WNTR24-cua (2.4 kpc), the cumulative Galactic extinction is mag, rising to only mag along the full line of sight, indicating that the source already lies behind most of the Galactic dust in this direction. In contrast, WNTR24-egv (2.52 kpc) shows mag out to its distance compared to mag for the entire column, implying that only part of the Galactic dust lies in the foreground and a substantial fraction remains behind the source. The much higher extinction estimate from the SED fit ( mag) therefore indicates significant local extinction in the circumstellar environment. For the analysis in Section 3.4, we adopt the extinction estimates from the SED fits.
@@
3.4 Luminosities and mass accretion rates
We estimate the mass accretion rates of both FUor outbursts by combining physical parameters derived from pre-main-sequence stellar models with bolometric luminosities obtained from the SEDs. For each system, we integrate the extinction-corrected SED between 1 and 30 m to determine the total luminosity before and after outburst (following Gordon et al., 2023). The post-outburst photometry used in this analysis was obtained in 2024, years after the outburst onset for each event.
For WNTR24-cua, we adopt mag and find a pre-outburst luminosity of and post-outburst luminosity of . In WNTR24-egv, where mag, we estimate and .

### v2

@@ Abstract
I Introduction
Young stellar objects (YSOs) grow through episodic accretion, where material from a surrounding disk falls onto the star in bursts rather than at a steady rate (Fischer et al., 2023, and references therein). These accretion outbursts significantly increase the system’s luminosity and play a major role in stellar mass assembly (Kenyon et al., 1990; Hartmann and Kenyon, 1996; Vorobyov and Basu, 2005). Episodic accretion has been proposed as a solution to the “luminosity problem,” in which protostars appear dimmer than expected if they were gaining mass at a constant rate throughout their evolution (Kenyon et al., 1990; Dunham and Vorobyov, 2012). In addition to influencing how stars gain mass, these bursts can alter the thermal and chemical structure of the disk, affecting processes such as snowline migration and early planet formation (Nayakshin et al., 2023; Lee et al., 2019; Jørgensen et al., 2013; Kóspál et al., 2023). Discovering and characterizing these outbursts remains essential for understanding their physical origins and role in star formation.
The most extreme form of episodic accretion in YSOs is FU Orionis-type (FUor) outbursts, characterized by large luminosity increases (mag 4–6) and lasting decades or more (Hartmann and Kenyon, 1996; Audard et al., 2014; Fischer et al., 2023). FUor outbursts sustain accretion rates of up to , contributing significantly to stellar mass growth in the earliest stages of evolution (Hartmann and Kenyon, 1996; Vorobyov and Basu, 2015). In contrast, another common category of bursts, EX Lupi-type outbursts (EXors), are less luminous (mag 2–3), shorter-lived (weeks to a few years), and contribute less significantly to the stellar mass budget (Ashraf et al., 2024; Cruz-Sáenz de Miera et al., 2023). Despite the proposed significance of FUors in mass accumulation, they are observationally rare. The classic compilation of Connelley and Reipurth (2018) reported fewer than 35 FUor-type events had been confirmed using spectroscopic techniques, albeit with most objects having sparse or no photometric coverage of the outburst.
@@ I Introduction
Recent wide-field IR surveys have proven essential for uncovering outbursts in heavily embedded YSOs, including classes of FUor outbursts previously missed in optical monitoring. The VISTA Variables in the Via Lactea (VVV; Minniti et al. 2010) and its extension VVVX enabled the discovery of dozens of high-amplitude eruptive YSOs, including several FUor candidates in the Galactic plane (Contreras Peña et al., 2017b, a, 2024; Guo et al., 2024a). Similarly, mid-infrared (MIR) monitoring with WISE (Wright et al., 2010) and NEOWISE (Mainzer et al., 2014) uncovered a growing population of deeply embedded eruptive variables (Hillenbrand et al., 2018; Ashraf et al., 2024; Tran et al., 2024; Contreras Peña et al., 2025). While the VVV and NEOWISE searches were largely implemented via archival photometric searches together with spectroscopic follow-up many years after the outburst onset, the advent of real-time transient discovery techniques is enabling prompt identification and early spectroscopic follow-up of these outbursts. The ongoing Palomar Gattini-IR (PGIR) survey has uncovered several new NIR-bright FUor events in the northern hemisphere (Hillenbrand et al., 2021, 2025) and continues to provide a key source of detections, especially toward the Galactic plane (Earley et al., in prep.).
@@
In addition to discovery, multi-color infrared photometry and broad-band spectroscopy play a key role in confirming YSO outbursts and enabling detailed studies of accretion disk evolution. These observations also identify and exclude interlopers, including Mira variables, cataclysmic variables, and classical novae. Multi-wavelength monitoring probes accretion disk formation via changes in NIR and MIR colors—where delays between shorter and longer wavelength bands are critical to understanding the outburst triggering mechanism and instability propagation (outside-in or inside-out; Bell et al. 1995; Cleaver et al. 2023; Nayakshin and Elbakyan 2024). Simultaneously, such coverage traces the development of hot inner disk regions, the clearing or rebuilding of circumstellar material and the presence of infalling envelopes or outflows (Hillenbrand and Rodriguez, 2022; Caratti o Garatti et al., 2017). Connelley and Reipurth (2018) demonstrate that NIR spectroscopy provides a robust basis for classifying FUor-type objects, which exhibit a distinctive combination of strong molecular absorption, weak or absent emission lines, low-gravity spectral features, and a mixed-temperature spectrum that becomes progressively cooler at longer wavelengths.
@@
The Wide-Field Infrared Transient Explorer (WINTER) is a new NIR time-domain survey on a dedicated 1-meter robotic telescope at Palomar Observatory (Lourie et al., 2020; Frostig et al., 2024, 2025). Each pointing covers a 1.2 field of view with a 1.1” pixel scale and the option of observing in Y, J, or a shortened H (Hs) filters, centered on 1.02, 1.25, and 1.60 . WINTER covers the Galactic plane during the summer months, running multiple surveys at an approximately two-week cadence. Unlike VVV and PGIR which almost entirely operated their time-domain surveys in a single filter (in Ks and J band, respectively), WINTER carries out multi-color cadenced surveys of the Galactic plane. Like PGIR, WINTER also includes the infrastructure for rapid discovery of transients with a real-time image subtraction pipeline, operating with smaller field of view (1.2 vs. 25 ) but substantially higher depth (median vs. AB mag) and spatial resolution (1.1″vs 8.7″) than PGIR (De et al., 2020).
@@
We present the observations of WNTR24-cua and WNTR24-egv in WINTER, WISE, and other surveys, the source environment, and spectroscopic follow-up in Section II. In Section III, we analyze the observations, presenting a study of the host environments, the progenitors of the outbursts, and spectroscopically classify the FUors. We discuss the implications for future observations of YSO outbursts with WINTER and conclude in Section IV.
II Observations
@@
II.1 Discovery, NEOWISE, and WINTER initial observations
@@ II.2 Optical Observations
II.3 Light curve evolution
@@
WNTR24-cua was not detected in WISE until early 2016, when it began a sharp rise in the band, brightening by 1.7 magnitudes over 364 days (Figure 2). This was followed by a decline of 0.5 magnitudes in 2019, after which the source exhibited a more gradual increase of 0.62 magnitudes between 2020 and 2024. The initial rise was not seen in , likely due to the band’s 1 magnitude shallower depth; however, near peak brightness the source was detected at mag, with a color of mag. By 2024, the emission became bluer with mag. The brightening is also evident in VVV , which shows the onset of the outburst in 2016, with the progenitor detected prior to the rise in , , and (data from Guo et al. 2024a). In contrast, Gaia -band observations do not detect the source up through May 2017, but then observe the outburst three months later, implying a rapid optical rise after the start of the NIR and MIR outburst.
@@
II.4 Archival imaging of the source environment
@@ III.1 Environment
WNTR24-egv lies in close proximity to the H ii region [DLW84] G35.2S, with G35.2N also nearby. Using Gaia EDR3 parallax measurements ( mas; Gaia Collaboration 2020), we estimate a distance of kpc to the region. The projected angular separation between WNTR24-egv and the center of G35.2S is , corresponding to a physical offset of only pc at this distance, indicating a likely association. A dense molecular core (W48 89) identified by Keown et al. (2019) is essentially identical in position with WNTR24-egv. Many sources in the immediate vicinity were identified by Kuhn et al. (2021) as YSOs, although not the progenitor of WNTR24-egv.
@@
Following the procedure outlined in Tran et al. 2024, we also estimate the distance to the outbursts by identifying nearby candidate YSOs with known distances. First, we take sources within 20’ of the FUors from the Spitzer (SEIP) source list (SSTSL2, Spitzer Science Center (SSC) and Infrared Science Archive (IRSA), 2021) and the 2MASS All Sky Catalog of point sources (Cutri et al., 2003) and apply a color cut of and to identify sources with IR-color excess as likely YSOs. IR-excess contaminants, such as star-forming galaxies, are eliminated following the procedures outlined in Gutermuth et al. (2009).
@@
This results in 466 and 273 candidate YSOs for WNTR24-cua and WNTR24-egv respectively. For WNTR24-cua, 17 of the candidate YSOs have counterparts in the Gaia DR3 catalog (Gaia Collaboration, 2020) with measured parallaxes, yielding a median distance of kpc. Due to the discrepancies in the estimated distances presented in Guo et al. 2024a, we use the estimate of 2.4 kpc for the analyses in Section III.3. Similarly, for WNTR24-egv, 10 of the candidate YSOs have Gaia distances ranging from 0.2 to 2.2 kpc. However, their relatively uniform spatial distribution offers no clear basis for a reliable distance estimate, and many of the stars in this region are likely too obscured for Gaia detections. We therefore adopt the distance of the nearby H ii region at 2.52 kpc.
III.2 Spectroscopic features and classification
@@
Both WNTR24-cua and WNTR24-egv are spectroscopically similar to other FUors at NIR wavelengths (as explored in detail in Connelley and Reipurth, 2018). The spectra are shown in Figure 3 with another FUor, Gaia17bpi (Hillenbrand et al., 2018), for comparison. All three spectra demonstrate a “mixed-temperature” spectrum characteristic of FUors. At the shortest infrared wavelengths (around 1 m), the spectra resemble a late-F to early-G type star at T6000 K. However, at longer wavelengths the spectra are consistent with an M-type star at T3000 K with deep molecular absorption features. For example, there is prominent absorption from the 12CO band-head (around ) and H2O band (m and m). Molecular absorption from VO and TiO is strongest in Gaia17bpi, present but weaker in WNTR24-cua, and not confidently detected in WNTR24-egv. Additionally, there are atomic lines from Al I, Ca I, Mg I, Na I, and Si I, along with H Paschen lines in both objects.
@@
Both of the outbursts show strong He I 10830 blue-shifted absorption, which is evidence for a strong wind (Connelley and Reipurth, 2018). We estimate a FWHM velocity of approximately 231 km/s and 305 km/s for WNTR24-cua and WNTR24-egv, respectively, and corrected for instrumental broadening. Measuring against the blue end of the absorption feature, we estimate a maximum velocity of 221 km/s and 321 km/s, for WNTR24-cua and WNTR24-egv.
@@
Based on these spectroscopic properties, we classify WNTR24-cua and WNTR24-egv as FUors following the criteria outlined in Connelley and Reipurth (2018). We measure the equivalent widths of CO, Na I, and Ca I, as outlined in Messineo et al. (2021), and compare to bona fide FUors from the Connelley and Reipurth (2018) sample (Figure 4). Both WNTR24-cua and WNTR24-egv fall within the bona fide sample, with notably large CO equivalent widths. Additionally, both sources exhibit the eight NIR spectral characteristics of FUors outlined in Connelley and Reipurth (2018). These characteristics include CO absorption and molecular band absorption, metallic, hydrogen, and He I lines, and changing of spectral type with wavelength.
III.3 Spectral energy distribution
@@
To place the SED-based extinction estimates in context, we also derive values from spectral dereddening and Galactic dust maps. First, we estimate the visual extinction () toward each source by dereddening their near-infrared spectra using the Cardelli et al. (1989) extinction law () until their continuum slopes match that of Gaia 17bpi, which has a known extinction of mag. This comparison yields relative extinctions of mag and mag, corresponding to values of mag for WNTR24-cua and mag for WNTR24-egv. We then compared these values to predictions from the Bayestar 3D Galactic dust map (Green et al., 2019). For WNTR24-cua (2.4 kpc), the cumulative Galactic extinction is mag, rising to only mag along the full line of sight, indicating that the source already lies behind most of the Galactic dust in this direction. In contrast, WNTR24-egv (2.52 kpc) shows mag out to its distance compared to mag for the entire column, implying that only part of the Galactic dust lies in the foreground and a substantial fraction remains behind the source. The much higher extinction estimate from the SED fit ( mag) therefore indicates significant local extinction in the circumstellar environment. For the analysis in Section III.4, we adopt the extinction estimates from the SED fits.
@@
Figure 5 also presents the dereddened pre-outburst SEDs assuming mag for WNTR24-cua and mag for WNTR24-egv (following extinction curves from Gordon et al. 2023, which builds upon Gordon et al. 2021; Decleir et al. 2022). These values represent estimates of the total foreground extinction, which likely includes both (mostly) interstellar and circumstellar components that are difficult to separate. The figure also presents one dereddened, representative photospheric model to show a simplified estimate of the intrinsic luminosity of just the pre-outburst object without the dust modeled by black bodies.
III.4 Luminosities and mass accretion rates
We estimate the mass accretion rates of both FUor outbursts by combining physical parameters derived from pre-main-sequence stellar models with bolometric luminosities obtained from the SEDs. For a simplified estimate of the total luminosity, we integrate the extinction-corrected SED between 1 and 30 m for both the pre- and post-outburst luminosity ( and ), which includes stellar, disk, and envelope emission. This provides a consistent measure of the systems’ total radiative output, including both reprocessed and accretion-powered components, though it ignores the potentially complex geometry of the system and likely overestimates the intrinsic luminosity of the central source due to the inclusion of reprocessed emission from the surrounding envelope. To isolate just the luminosity from the proto-star (), we integrate the extinction-corrected () photosphere model to isolate the central source from the surrounding dust. As a note, the post-outburst photometry used in this analysis was obtained in 2024, years after the outburst onset for each event.
For WNTR24-cua, we adopt mag and find a total pre-outburst luminosity of , and post-outburst luminosity of . In WNTR24-egv, where mag, we estimate and .